Find the solution to the following System of Linear Equations
y=4x-2
y=-2x+4

Answer by ikleyn(52790)

.

Left sides are identical --- hence, right side are equal.


It gives you an equation in one single unknown x


    4x - 2 = -2x + 4


Simplify and find x


    4x + 2x = 4 + 2

       6x   =   6

        x    = 6/6 = 1.


The from first given equation 


    y = 4*1 - 2 = 2.


Answer.  x= 1;  y= 2.

Solved.

MEMORIZE this mantra: it is the shortest way to get the solution/(the answer).